# A Reduction-Driven Local Search for the Generalized Independent Set Problem

### v1

@@ Abstract
The Generalized Independent Set (GIS) problem extends the classical maximum independent set problem by incorporating profits for vertices and penalties for edges. This generalized problem has been identified in diverse applications in fields such as forest harvest planning, competitive facility location, social network analysis, and even machine learning. However, solving the GIS problem in large-scale, real-world networks remains computationally challenging. In this paper, we explore data reduction techniques to address this challenge. We first propose 14 reduction rules that can reduce the input graph with rigorous optimality guarantees. We then present a reduction-driven local search (RLS) algorithm that integrates these reduction rules into the pre-processing, the initial solution generation, and the local search components in a computationally efficient way. The RLS is empirically evaluated on 278 graphs arising from different application scenarios. The results indicates that the RLS is highly competitive – For most graphs, it achieves significantly superior solutions compared to other known solvers, and it effectively provides solutions for graphs exceeding 260 million edges, a task at which every other known method fails. Analysis also reveals that the data reduction plays a key role in achieving such a competitive performance.
@@ 1 Introduction
Graphs are essential tools for modeling complex systems including social networks, geometric networks, and forest management systems. The analysis of graph data has led to the development of numerous combinatorial problems that help address real-world challenges. In this paper, we study the generalized independent set (GIS) problem. Let $G=(V,E=E_{p}\cup E_{r})$ be an undirected graph where $V$ is the vertex set and $E$ is the edge set that is partitioned into a permanent edge set $E_{p}$ and a removable edge set $E_{r}$. Each vertex $v\in V$ is assigned a profit $w(v)\in\mathbb{R}$. Each removable edge $e\in E_{r}$ is associated with a penalty $p(e)\in\mathbb{R}$. The generalized independent set of the graph $G$ is a vertex set $I\subseteq V$ such that no permanent edge exists between any two vertices of $I$ in $G$. Then, the GIS problem asks for a generalized independent set $I$ of maximum net benefit, which is the total vertex profits of $I$ minus the total penalties of removable edges whose end vertices are in $I$. An example is shown in Figure 1.
Figure 1: A toy example of the GIS problem. The set of red vertices is a generalized independent set of net benefit of $2-1+6+5=12$.
@@
When all the edges are permanent edges, the GIS problem equals the well-known maximum independent set problem (MIS) which asks for a maximum set of non-adjacent vertices from a given graph. As the GIS allows a more flexible definition of vertex profit and edge penalties, it captures a broader range of real-world scenarios than the classical independent set problem. We leave detailed examples of its applicability in diverse domains in Section 2.1. However, since MIS is NP-hard, W[1]-hard and difficult to approximate (Cygan et al. 2015), the equivalence indicates that the GIS is also computationally hard.
@@
Existing works have proposed various approaches to solve the GIS problem practically. Exact solution methods include mixed integer linear programming (Colombi et al. 2017) and unconstrained binary quadratic programming (Hosseinian and Butenko 2019). They can solve instances with up to 400 vertices and 71,820 edges in three hours so far (Zheng et al. 2024). On the other hand, heuristic search methods have demonstrated their effectiveness in providing high-quality solutions within reasonable computational time (e.g., Colombi et al. (2017), Nogueira et al. (2021)). For example, the state-of-the-art ALS algorithm proposed by Zheng et al. (2024) achieves near-optimal solutions on all standard benchmark instances and scales to problems with up to 18,000 vertices. Nevertheless, we observe that these heuristic algorithms struggle to maintain solution quality in even larger instances.
@@
The data reduction, which reduces input size while preserving problem optimality, has recently emerged as a promising technique for improving algorithm scalability on large instances. The technique has been successfully applied to problems such as independent set, vertex cover, and maximum clique (Großmann et al. 2023, Abu-Khzam et al. 2022, Verma et al. 2015, Walteros and Buchanan 2020, Lamm et al. 2019). For the GIS problem, research on data reduction is still in its early stages. There are three main challenges in developing data reduction techniques for the GIS problem: (1) The complex interplay among vertex profits, edge penalties, and the underlying graph structure, (2) Translating reduction rules into efficient, practical algorithms, often requiring incremental implementations (Akiba and Iwata 2016, Chang et al. 2017), (3) Integrating reductions into local search algorithms while balancing search time and efficiency.
@@
    We propose an efficient reduction-driven local search algorithm, named RLS, to solve the GIS problem. We incorporate our proposed reduction rules into three main components of RLS: the pre-processing, which incrementally reduces the size of original input instance, the initial solution generation, which interplays the vertex selection and data reduction, and the neighborhood search, which uses a novel reduction-based move operator to guide the search trajectory.
@@
    We evaluate the proposed RLS algorithm on a total of 278 instances. On small instances, RLS consistently achieves the known optimal solutions with significantly shorter running time than exact algorithms. On large instances, RLS exhibits superior scalability over existing heuristic algorithms in solution quality. Notably, RLS successfully handles instances with over 260 million edges, where existing heuristic methods fail. Our experimental results also demonstrate that data reduction techniques significantly improve the computational efficiency of RLS.
The rest of the paper is organized as follows. Section 2 reviews applications and related works, followed by Section 3 which introduces the notations and definitions. In Section 4, we present our reduction rules for GIS. Then, in Section 5, we describe the RLS algorithm. Section 6 presents computational results and comparisons with existing methods. In the last section, we conclude the paper and outline future directions.
@@ 2.1 Applications Related to GIS
Forest harvest planning. The planning of forest harvest involves selecting forest areas (cells) for timber, aiming to maximize benefits like timber value while minimizing ecological damage from excessive cutting. As a GIS problem, forest cells are vertices weighted by harvesting benefits, and edges represent penalties for harvesting adjacent cells (Hochbaum and Pathria (1997)).
Competitive facility location. In determining optimal locations for service facilities, each potential site has a utility value. However, the utility diminishes when multiple facilities compete for customers in the same area. The facility location problem, modeled with vertices as facilities and edges as geographic relationships, can be framed as a GIS problem to maximize utilities (Hochbaum (2004)). Similar spatial conflicts appear in airplane seating (Pavlik et al. 2021) and label placement (Mauri et al. 2010).
Social network analysis. The GIS problem has broad applications in social network analysis. A maximum independent set corresponds to the complement of a minimum vertex cover and is equivalent to a maximum clique in the graph’s complement (Sanchis and Jagota (1996)). These properties make GIS useful for tasks such as detecting cliques or communities (Gschwind et al. (2021)) and optimizing coverage and reach within social networks (Puthal et al. (2015)).
@@ 2.2 Existing Algorithms
Exact Algorithms. In pursuit of optimal solutions, Hochbaum (2004) introduced an integer programming model that considers both nodes and edges as variables, with each edge associated with a specific constraint. However, this model relies on binary variables for removable edges, resulting in large formulations; for example, a graph with 400 vertices requires 80,000 variables, making direct linear programming infeasible. To address this, Colombi et al. (2017) reformulated the GIS problem as an unconstrained binary quadratic program (UBQP), and developed a branch-and-cut algorithm, along with additional valid inequalities. Note that they also generated a total of 216 random instances which often serve as benchmark datasets in the later work. Hosseinian and Butenko (2019) employed another linear program to bound the optimal solution, which is later incorporated into a branch-and-bound framework. Recently, Zheng et al. (2024) used a Lagrangian relaxation in their branch-and-bound framework to estimate tight upper bounds. They showed that their algorithm is able to solve the GIS instances with up to 400 vertices within $3$ hours.
@@
Heuristic Algorithms. Kochenberger et al. (2007) also developed a UBQP formulation for the GIS problem solved via tabu search. Colombi et al. (2017) enhanced performance by integrating advanced tabu search with UBQP. Nogueira et al. (2021) directly solved the original problem by using an iterated local search heuristic. They explored move operations of adding one or two vertices to the incumbent solution. More recently, Zheng et al. (2024) proposed a more refined local search heuristic, namely ALS, which explores three move operations – adding one vertex to the solution, dropping one vertex from the solution, and replacing a vertex in the solution with another one not in the solution. The ALS algorithm reported state-of-the-art results on the set of 216 benchmark instances proposed by Colombi et al. (2017). Their algorithm is able to find high-quality solutions on large instances with up to $17903$ vertices within $5$ minutes. However, its performance on even larger graphs, such as the uk2002 graph with 18 million vertices, remains unknown.
@@
Data Reduction Techniques. The data reduction techniques are originally introduced in the framework of parameterized complexity analysis by the name of kernelization (Cygan et al. 2015). Recently, data reduction techniques are intensively used for practically tackling MIS (Abu-Khzam et al. 2022). For example, Akiba and Iwata (2016) shows that branch-and-reduce outperforms pure branch-and-bound, while Dahlum et al. (2016) and Chang et al. (2017) combine reductions with local search and heuristics to achieve near-optimal performance. These techniques enable handling graphs with millions of vertices and edges (Lamm et al. 2019, Xiao et al. 2021). Besides, the effectiveness of reductions has also been acknowledged in other fundamental problems like the clique problem (Walteros and Buchanan 2020), and the matching problem (Koana et al. 2021). However, as for the GIS problem, no existing work focuses on reduction rules. And extending existing reductions to vertex- and edge-weighted graphs for the GIS problem remains challenging.
@@ 3 Preliminary
By the definition, the input instance of the GIS problem is an (undirected weighted) graph $G=(V,E=E_{p}\cup E_{r})$ where each vertex $v\in V$ is associated with a profit $w(v)\in\mathbb{R}$ and each removable edge $e\in E_{r}$ is associated with a penalty $p(e)$. In order to reformulate the problem in a uniform way, we extend the definition by associating each edge $e\in E_{p}$ also with a penalty $p(e)=M$ where $M\in R$ is a sufficiently large value, e.g., the absolute sum of all vertex profits and removable edge penalties, $\sum_{v\in V}|w(v)|+\sum_{e\in E_{r}}|p(e)|$). For a set $S\subseteq V$, we formally define $nb(S)$ as the net benefit of $S$, i.e.
$nb(S)=\sum_{v\in S}w(v)-\sum_{\{u,v\}\subseteq S,(u,v)\in E}p(u,v).$
Then, the GIS problem is equal to finding a set $S\subseteq V$ that maximizes the net benefit of $S$ from a given graph $G$. The optimal objective value of the GIS problem in $G$ is denoted by $\alpha(G)$, i.e.
$\alpha(G)=\max_{S\subseteq V}nb(S)$
We also denote $MGIS(G)$ as the set of all vertex sets $S$ with optimal objective value. It is clear that every $S\in MGIS(S)$ is a generalized independent set which is defined in the introductory section. We say a generalized independent set $S\subseteq V$ is maximal if there is no other generalized independent set $S^{\prime}$ such that $S\subset S^{\prime}$ and $nb(S^{\prime})\geq nb(S)$.
For convenience, we import more notations in Table 1. Note that we use $\tilde{w}(v)$ to represent the profit of $v$ plus the non-negative penalties of edges that are incident to $v$. The $w^{+}(S)$ is an upper bound on the net benefit of any subset of $S$. For example, in Figure 1, $\tilde{w}(u)=2+1=3$. When $S$ denotes the set of red vertices in Figure 1, $w^{+}(S)=3+6+5=14$.
In the rest of the paper, we use notation $\coloneqq$ to represent assignment and $=$ to indicate equivalence.
@@
$G=(V,E)$ a graph $G$ with vertex set $V$ and edge set $E$
$E_{p},E_{r}$ the set of pernament edges and the set of removable edges respectively
$w(v),p(u,v)$ the profit of including a vertex $v$, and the penalty of including both vertices $u$ and $v$
$N(v),N_{p}(v),N_{r}(v)$ the sets of all neighbors, permanent neighbors and removable neighbors of $v$, respectively
@@
the sets of all neighbors, permanent neighbors and removable neighbors of $v$ including $v$
itself, respectively
@@
$\tilde{w}(v)=w(v)+\sum_{x\in N_{r}(v)}\max(0,-p(v,x))$
$S,G\setminus S$ a vertex subset of $G$, and the graph obtained by removing $S$ and all incident edges
@@
$nb(S)$ the net benefit of vertex set $S$, $nb(S)=\sum_{v\in S}w(v)-\sum_{\{u,v\}\subseteq S}p(u,v)$
$\alpha(G),MGIS(G)$ the largest net benefit value, and the set of all vertex sets whose net benefit is $\alpha(G)$
Integer Programming. The GIS problem can be easily formulated as (Colombi et al. 2017):
$\alpha(G)=\max\quad$ $\sum_{i\in V}w(i)x_{i}-\sum_{\{i,j\}\in E_{r}}p(i,j)y_{ij}$ (1)
s.t. $x_{i}+x_{j}\leq 1$ $(i,j)\in E_{p}$ (2)
$x_{i}+x_{j}-y_{ij}\leq 1$ $(i,j)\in E_{r}$ (3)
@@
$y_{ij}\in\{0,1\}$ $(i,j)\in E_{r}$ (5)
where binary variables $x_{i}$ indicate whether node $i$ is selected, and binary variables $y_{ij}$ indicate whether the edge is included.

### v2

@@ Abstract
The Generalized Independent Set (GIS) problem extends the classical maximum independent set problem by incorporating profits for vertices and penalties for edges. This generalized problem has been identified in diverse applications in fields such as forest harvesting, competitive facility location, social network analysis, and even machine learning. However, solving the GIS problem in large-scale, real-world networks remains computationally challenging. In this paper, we explore data reduction techniques to address this challenge. We first propose 14 reduction rules that can reduce the input graph with rigorous optimality guarantees. We then present a reduction-driven local search (RLS) algorithm that integrates these reduction rules into the pre-processing, the initial solution generation, and the local search components in a computationally efficient way. The RLS is empirically evaluated on 278 graphs drawn from different application scenarios. The results indicate that the RLS is highly competitive – For most graphs, it achieves significantly superior solutions compared to other known solvers, and it effectively provides solutions for graphs exceeding 260 million edges, a task at which every other known method fails. Analysis also reveals that data reduction plays a key role in achieving such a competitive performance.
@@ 1 Introduction
Graphs are essential tools for modeling complex systems including social networks, geometric networks, and forest management systems. The analysis of graph data has led to the development of numerous combinatorial problems that help address real-world challenges. In this paper, we study the Generalized Independent Set (GIS) problem. Let $G=(V,E=E_{p}\cup E_{r})$ be an undirected graph where $V$ is the vertex set and $E$ is the edge set that is partitioned into a permanent edge set $E_{p}$ and a removable edge set $E_{r}$. Each vertex $v\in V$ is assigned a profit $w(v)\in\mathbb{R}$. Each removable edge $e\in E_{r}$ is associated with a penalty $p(e)\in\mathbb{R}$. A generalized independent set of the graph $G$ is a vertex set $I\subseteq V$ such that no permanent edge exists between any two vertices of $I$ in $G$. Then, the GIS problem asks for a generalized independent set $I$ of maximum net benefit, which is the total vertex profits of $I$ minus the total penalties of removable edges whose end vertices are in $I$. An example is shown in Figure 1.
Figure 1: A toy example of the GIS problem. The set of red vertices is a generalized independent set of net benefit of (2+6+5)1=12(2+6+5)-1=12.
@@
When all the edges are permanent and each vertex has a profit of 1, the GIS problem becomes the well-known maximum independent set problem (MIS), which asks for a maximum-cardinality set of non-adjacent vertices in a given graph. As the GIS problem allows a more flexible definition of vertex profit and edge penalties, it captures a broader range of real-world scenarios than the classical independent set problem. We give detailed examples of applications in diverse domains in Section 2.1. However, since MIS is NP-hard, W[1]-hard and difficult to approximate (Cygan et al. 2015), the equivalence above indicates that the GIS problem is also computationally hard.
@@
Existing works have proposed various approaches to solve the GIS problem practically. Exact solution methods include mixed integer linear programming (Colombi et al. 2017) and unconstrained binary quadratic programming (Hosseinian and Butenko 2019). They can solve instances with up to 400 vertices and 71,820 edges in three hours so far (Zheng et al. 2024). On the other hand, heuristic search methods have demonstrated their effectiveness in providing high-quality solutions within reasonable computational time (e.g., Colombi et al. (2017), Nogueira et al. (2021)). For example, the most recent algorithm, named Adaptive Local Search (ALS), proposed by Zheng et al. (2024) achieves near-optimal solutions on all standard benchmark instances and scales to problems with up to 18,000 vertices. Nevertheless, we observed that these heuristic algorithms struggle to maintain solution quality in even larger instances.
@@
Data reduction, which reduces input size while preserving problem optimality, has recently emerged as a promising technique for improving algorithm scalability on large instances. The technique has been successfully applied to problems such as independent set, vertex cover, and maximum clique (Großmann et al. 2023, Abu-Khzam et al. 2022, Verma et al. 2015, Walteros and Buchanan 2020, Lamm et al. 2019). For the GIS problem, research on data reduction is still at its early stages. There are three main challenges in developing data reduction techniques for the GIS problem: (1) The complex interplay among vertex profits, edge penalties, and the underlying graph structure, (2) Translating reduction rules into efficient, practical algorithms, which often requires incrementally applying the reduction rules (Akiba and Iwata 2016, Chang et al. 2017), (3) Integrating reductions into local search algorithms while balancing search time and efficiency.
@@
    We propose an efficient reduction-driven local search algorithm, named RLS, to solve the GIS problem. We incorporate our proposed reduction rules into three main components of the RLS: the pre-processing, which incrementally reduces the size of the original input instance, the initial solution generation, which interplays the vertex selection and data reduction, and the neighborhood search, which uses a novel reduction-based move operator to guide the search trajectory.
@@
    We evaluate the proposed the RLS algorithm on a total of 278 instances. On small instances with up to 400 vertices and 80,000 edges, the RLS consistently achieves the known optimal solutions with significantly shorter running time than exact algorithms. On large instances, the RLS exhibits superior scalability over existing heuristic algorithms in solution quality. Notably, the RLS successfully handles instances with over 18 million vertices and over 260 million edges, where existing heuristic methods fail. Our experimental results also demonstrate that data reduction techniques significantly improve the computational efficiency of the RLS.
The rest of the paper is organized as follows: Section 2 reviews applications and related works, followed by Section 3 which introduces the notations and definitions. In Section 4, we present our reduction rules for the GIS problem. Then, in Section 5, we describe the RLS algorithm. Section 6 presents computational results and comparisons with existing methods. In the last section, we conclude the paper and outline future directions.
@@ 2 Background
In this section, we provide some application examples of the GIS model and review existing GIS algorithms.
@@ 2.1 Applications Related to GIS
Forest harvesting. The planning of forest harvesting involves selecting forest areas (cells) for timber, aiming to maximize benefits like timber value while minimizing ecological damage from excessive cutting. As a GIS problem, forest cells are vertices weighted by harvesting benefits, and edges represent penalties for harvesting adjacent cells (Hochbaum and Pathria (1997)).
Competitive facility location. In determining optimal locations for service facilities, each potential site has a utility value. However, the utility diminishes when multiple facilities compete for customers in the same area. The facility location problem, modeled with vertices as facilities and edges as geographic relationships, can be framed as a GIS problem to maximize utilities (Hochbaum (2004)). Similar spatial conflicts exist in airplane seating (Pavlik et al. 2021) and label placement (Mauri et al. 2010).
Social network analysis. Potentially, the GIS problem has broad applications in social network analysis. A maximum independent set corresponds to the complement of a minimum vertex cover and is equivalent to a maximum clique in the graph’s complement (Sanchis and Jagota (1996)). These properties make GIS useful for tasks such as detecting cliques or communities (Gschwind et al. (2021)) and optimizing coverage and reach within social networks (Puthal et al. (2015)).
@@ 2.2 Existing Algorithms
Exact Algorithms. In pursuit of optimal solutions, Hochbaum (2004) introduced an integer linear programming model that assigns variables to each node and edge, and associates each edge with a specific constraint. As a result, the integer linear program for a graph of 2000 vertices (e.g. C2000.5 in the DIMACS competition set) has approximately 500,000 variables and 500,000 constraints, which is quite large for modern ILP solvers like CPLEX. (It fails to even report a solution within 3 hours for this instance.) Then, Colombi et al. (2017) reformulated the GIS problem as an unconstrained binary quadratic program (UBQP), and developed a branch-and-cut algorithm, along with additional valid inequalities. Note that they also generated a total of 216 random instances which often serve as benchmark datasets in the later work. Hosseinian and Butenko (2019) employed another quadratic program to bound the optimal solution, which was later incorporated into a branch-and-bound framework. Recently, Zheng et al. (2024) used a Lagrangian relaxation in their branch-and-bound framework to estimate tight upper bounds. Their experiments showed that their algorithm is able to solve the GIS instances with up to 400 vertices within $3$ hours.
@@
Heuristic Algorithms. Kochenberger et al. (2007) also developed a UBQP formulation of the GIS problem and solved it using tabu search. Colombi et al. (2017) enhanced performance by integrating advanced tabu search with UBQP. Nogueira et al. (2021) directly solved the original problem by using an iterated local search heuristic. They explored move operations of adding one or two vertices to the incumbent solution. More recently, Zheng et al. (2024) proposed a more refined local search heuristic“AL” to explores three move operations – adding one vertex to the solution, dropping one vertex from the solution, and replacing a vertex in the solution with another one not in the solution. The ALS algorithm has been reported to achieve state-of-the-art performance on the standard benchmark set of 216 instances proposed by Colombi et al. (2017). In particular, it is able to find high-quality solutions for instances with up to $17,903$ vertices within $5$ minutes. However, its performance on graphs with more than 20,000 vertices has not been systematically evaluated. Our experiments show that ALS quickly becomes impractical on such large-scale instances.
@@
Data Reduction Techniques. Data reduction techniques for integer programming models can be traced back several decades, referred to as variable fixing and coefficient reduction (crowder1983solving). Then, the parameterized complexity community has also adopted this technique, known as kernelization, to achieve strong theoretical bounds on the time complexity of many hard combinatorial graph problems (Cygan et al. 2015). Recently, data reduction techniques are intensively used for practically tackling MIS (Abu-Khzam et al. 2022). For example, Akiba and Iwata (2016) shows that branch-and-reduce outperforms pure branch-and-bound, while Dahlum et al. (2016) and Chang et al. (2017) integrate reductions with heuristics to achieve near-optimal performance. These techniques enable handling graphs with millions of vertices and edges (Lamm et al. 2019, Xiao et al. 2021). Besides, the effectiveness of reductions has also been acknowledged in other fundamental problems like the clique problem (Walteros and Buchanan 2020), and the matching problem (Koana et al. 2021). However, as for the GIS problem, no existing work focuses on reduction rules. Extending existing reductions for the independent set problem to the GIS problem, where both vertices and edges are weighted, remains challenging.
@@ 3 Preliminary
According to the definition, an input instance of the GIS problem is an (undirected weighted) graph $G=(V,E=E_{p}\cup E_{r})$ where each vertex $v\in V$ is associated with a profit $w(v)\in\mathbb{R}$ and each removable edge $e\in E_{r}$ is associated with a penalty $p(e)\in\mathbb{R}$. For a vertex set $S\subseteq V$, we define $nb(S)$ as the net benefit of $S$, i.e.,
$nb(S)=\sum_{v\in S}w(v)-\sum_{u\in S,v\in S,e=\{u,v\}\in E_{r}}p(e).$
Then, a generalized independent set of $G$ is a vertex subset $S\subseteq V$ such that $G[S]$ contains no permanent edge. The GIS problem can be formulated as finding a generalized independent set such that its net benefit is maximized. We denote the optimal objective value of the GIS problem in $G$ as $\alpha(G)$, i.e.,
$\alpha(G)=\max_{S\subseteq V}nb(S)\quad s.t.\quad E_{p}\cap\{\{u,v\}\mid u,v\in S\}=\varnothing$
We also denote $MGIS(G)$ as the family of generalized independent sets $S\subseteq V$ that achieve the optimal objective value $\alpha(G)$. Each $S\in MGIS(G)$ corresponds to an optimal solution.
@@
$G=(V,E=E_{p}\cup E_{r})$
a graph $G$ with vertex set $V$ and edge set $E$,
$E_{p},E_{r}$ represent the set of permanent edges and the set of removable edges, respectively
$w(v),p(\{u,v\})$ the profit of including a vertex $v$, and the penalty of including both vertices $u$ and $v$
$N(v),N_{p}(v),N_{r}(v)$
subsets of $V$, representing the sets of vertices connected to $v$, vertices connected to $v$ by
a permanent edge and vertices connected to $v$ by a removable edge, respectively
@@
subsets of $V$, representing the sets $N(v)\cup\{v\}$, $N_{p}(v)\cup\{v\}$, $N_{r}(v)\cup\{v\}$, respectively
@@
$\tilde{w}(v)=w(v)+\sum_{x\in N_{r}(v)}\max(0,-p(\{v,x\}))$
S,G[S],E(G[S])S,{G[S],E(G[S])} a vertex subset of $G$, the subgraph induced by $S$, and the edge set of the induced subgraph
@@
$nb(S)$ the net benefit of a vertex set $S$, $nb(S)=\sum_{v\in S}w(v)-\sum_{u\in S,v\in S,e=\{u,v\}\in E_{r}}p(e)$
$\alpha(G),MGIS(G)$
the largest net benefit value achievable without including any permanent edge,
and the family of all subsets of $V$ attaining $\alpha(G)$ value under the same constraint
For convenience, we report more notations in Table 1. Note that we use $\tilde{w}(v)$ to represent the profit of $v$ plus the absolute sum of non-positive penalties of edges that are incident to $v$. Given that $S\subseteq V$ is a generalized independent set, $w^{+}(S)$ is an upper bound on the net benefit of $S$. For example, in Figure 1, $\tilde{w}(u)=2+2.5=4.5$. When $S$ denotes the set of red vertices in Figure 1, $w^{+}(S)=4.5+6+5=15.5$.
In the rest of the paper, we use the notation $\coloneqq$ to represent assignment and == to indicate equivalence.
Integer Programming. The GIS problem can be conveniently formulated as follows (Colombi et al. 2017):
$\alpha(G)=\max\quad$ $\sum_{i\in V}w(i)x_{i}-\sum_{\{i,j\}\in E_{r}}p(\{i,j\})y_{ij}$ (1)
s.t. $x_{i}+x_{j}\leq 1$ $\{i,j\}\in E_{p}$ (2)
$x_{i}+x_{j}-y_{ij}\leq 1$ $\{i,j\}\in E_{r}$ (3)
${x_{i}\geq y_{ij},x_{j}\geq y_{ij}}$ $\{i,j\}\in E_{r}$ (4)
@@
$y_{ij}\in\{0,1\}$ $\{i,j\}\in E_{r}$ (6)
where binary variables $x_{i}$ indicate whether the vertex $i$ is selected, and binary variables $y_{ij}$ indicate whether the removable edge $\{i,j\}$ is included.